OpenStudy (anonymous):

Calculate the surface area of a square pyramid if the side length of the base 4 cm and the slant height is 7 cm.

OpenStudy (luigi0210):

Do you know the formula for surface area?

OpenStudy (anonymous):

B +1/2ps

OpenStudy (luigi0210):

You sure?

OpenStudy (anonymous):

Yeap

jigglypuff314 (jigglypuff314):

http://math.about.com/od/formulas/ss/surfaceareavol_5.htm surface area of square pyramid = 2bs + b^2 plug in b=4 and s=7 so = 2(4)(7) + (4)^2 =